A cup of coffee (temperature = 190°F) is placed in a room whose temperature is 70°F. After five minutes, the temperature of the coffee has dropped to 160°F. How many more minutes must elapse before the temperature of the coffee is 130°F?
1
Expert's answer
2020-09-28T20:14:47-0400

As Newton's law of cooling is "T(t)=T_e+(T_0-T_e)e^{-kt}" where "T_e" is a temperature of environment, "T_0" is a temperature at time "t=0" . When

"k=-\frac{1}{t}\ln\frac{T_1-T_e}{T_0-T_e}" then "t_2=t_1(1-\ln{\frac{T_1-T_e}{T_2-T_e}}/{\ln\frac{T_1-T_e}{T_0-T_e}})" where "T_0=190^0F,T_1=160^0F,T_2=130^0F, t_1=5min" , then "t_2\approx12min"